### v2

# TimeFlow: Longitudinal Brain Image Registration and Aging Progression Analysis

Abstract
Predicting future brain states is crucial for understanding healthy aging and neurodegenerative diseases. Longitudinal brain MRI registration, a cornerstone for such analyses, has long been limited by its inability to forecast future developments, reliance on extensive dense longitudinal data, and the need to balance registration accuracy with temporal smoothness. In this work, we present TimeFlow, a novel framework for longitudinal brain MRI registration that overcomes all these challenges. TimeFlow leverages a U-Net architecture with temporal conditioning inspired by diffusion models, enabling accurate registration using only two images as input and facilitating prospective analyses through future image prediction. Unlike traditional methods, TimeFlow eliminates the demand for explicit smoothness regularizers and dense sequential data while maintaining temporal consistency and continuity. Experimental results highlight its superior performance in both future timepoint prediction and registration accuracy compared to state-of-the-art methods. Additionally, TimeFlow supports novel biological brain aging analyses, effectively differentiating neurodegenerative conditions from healthy aging, all without requiring segmentation thus avoiding non-trivial annotation and inconsistent segmentation flaws. This framework paves the way for accurate, data-efficient, and annotation-free prospective analyses of brain aging and chronic diseases.
Figure 1: Outstanding performance of TimeFlow for future brain MRI prediction. TimeFlow demonstrates the superior capability of future extrapolation compared to (i) NODER, which employs NeuralODE for geodesic regression, and (ii) VoxelMorph with linear scaling. Prediction errors in these methods are marked with red arrows. Given two previous timepoints (Baseline and Year 3), TimeFlow can extrapolate precise and smooth deformation to predict brain MRI in Year 6.
1 Introduction
Understanding how the human brain changes over time — both in the course of healthy aging and under the influence of neurodegenerative diseases — remains a central challenge in neurology and related fields [33, 34]. Such insights not only shed light on the underlying pathological processes but also pave the way for more individualized treatment planning and earlier, more targeted interventions [6]. Magnetic Resonance (MR) imaging has become the standard modality for capturing these longitudinal changes, owing to its high spatial resolution and excellent soft-tissue contrast. By repeatedly scanning the same individual, clinicians and researchers can visualize evolving patterns of atrophy, lesion progression, and other structural transformations as they unfold [21, 2].
Despite advances in longitudinal registration [17, 43], which aligns a subject’s brain images across multiple time points, two critical challenges persist: 1. Predicting future brain states and analyzing aging progression is one of the ultimate goals of longitudinal registration. While retrospective assessments of past changes are valuable, the ability to predict future brain transformations would dramatically enhance clinical utility [27] and enable clinicians to forecast atrophy trajectories and guide more effective interventions [23]. By extrapolating physically plausible transformations, longitudinal registration enables future scan prediction through simple image warping, avoiding the need for complex generative approaches. Moreover, comparing the predicted future brain changes with the actual observed brain state allows clinicians to assess whether the disease is advancing at an accelerated rate compared to historical progression. This insight can provide early warnings of disease exacerbation and help refine patient-specific prognostic models. Recent works [50, 3, 51] primarily focus on modeling the observed sequence of scans, allowing interpolation between observed timepoints but lacking the capability to extrapolate beyond the final scan, as illustrated in Fig. 1. This limitation makes it difficult to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly. 2. Ensuring temporal continuity with limited sequential data. Existing methods impose temporal smoothness constraints [49, 53, 9] and often require densely sampled longitudinal data (often more than five scans) [28, 36], which is impractical in typical clinical scenarios where only a few follow-up scans (<3) are available.
We introduce TimeFlow, a novel longitudinal brain MRI registration capable of predicting future brain changes. The key contribution is a novel constraint that ensures temporal consistency and continuity — removing the demand for smoothness regularizers. Moreover, TimeFlow can operate longitudinal registration with only paired images in inference, eliminating the requirement for long image sequences. Finally, TimeFlow can extrapolate to future time points and analyze aging progression without relying on segmentation masks. Our contributions are twofold:
  1. 1.
    Inspired by conditioning mechanisms used in diffusion models, TimeFlow employs a U-Net architecture conditioned on a temporal parameter reflecting biological aging. This temporal conditioning allows for generating temporally continuous and smooth deformation fields, even when only two images are used as input. Moreover, this approach enables future brain images prediction, surpassing state-of-the-art methods in accuracy and smoothness.
  2. 2.
    We develop a novel methodology for longitudinal brain aging and disease progression analysis, utilizing TimeFlow’s advanced predictive capabilities. Our approach quantifies relative aging progression using healthy subjects as a reference baseline, where biological aging is assumed to align with chronological aging. This method enables the identification and differentiation of subjects with Mild Cognitive Impairment (MCI) and Dementia, who typically exhibit accelerated biological aging compared to chronological aging. Remarkably, our method performs this analysis solely through registration, eliminating the need for brain structure annotations and avoiding potential segmentation-related error propagation.
-sho
2 Related Work
Cross-sectional registration
methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [41, 29], optical flow [35, 52], and diffeomorphic transformations [44, 47, 5]. However, they demand prohibitive computing time in the order of several hours to register one single subject, posing a significant barrier to clinical deployment. To address this, learning-based frameworks [4, 31, 37, 7, 14, 22] have emerged, dramatically reducing inference time and often improving registration accuracy. Nevertheless, all such methods rely on carefully tuned smoothness regularizers to ensure anatomically plausible deformation fields. An improper weighting factor of smoothness regularizers can either reduce fidelity or introduce unrealistic warps. Additional tuning or adaptive strategies are often needed to strike an optimal balance [18, 32, 48].
Longitudinal registration
methods [42, 19, 20, 11, 12, 36, 9, 24] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject. Two major strategies are often used to tackle this problem. First, an explicit smoothness constraint along the temporal dimension is incorporated to maintain coherent deformation fields over time [53, 28]. However, as with spatial smoothness, temporal regularization faces a similar trade-off: enforcing temporal continuity can inadvertently lower spatial accuracy. Second, many longitudinal methods require multiple, densely sampled scans (>10) to capture temporal dynamics [3, 50]. In real-world clinical scenarios, patients often have only a few follow-up scans ($\leq 3$), restricting model flexibility and practical implementation. Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies such as future image extrapolation and aging/disease progression risk prediction.
3 Methods
Figure 2: Left: Architecture of TimeFlow. TimeFlow $f_{\theta}$ integrates time $t$ information via adaptive instance normalization. The network takes the time $t$, the images at first $I_{0}$ and last visit $I_{L}$ as input, and predicts the temporal deformations. Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown. Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency. TimeFlow can predict bidirectional time-continuous deformation fields. The symmetric midpoint similarity forces the deformed images to be aligned at time $t_{k}$, while the midpoint interpolation flow loss encourages the deformation composition consistency. The extrapolation similarity and flow enforcement give information on what a realistic and reasonable extrapolation should be. Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$. The input images are highlighted by blue squares (Input), while the supervision are highlighted by dot-lined squares (Sup.).
3.1 Time-Continuous Registration
Given a sequence of longitudinal images $\{I_{k}\mid k=0,\ldots,L\}$, we aim to find a temporally continuous deformation trajectory that connects all the discrete time points. To achieve this, we construct a time-continuous registration function $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$. It takes the first visit $I_{0}$ and the last visit $I_{L}$ as input images and predicts the deformation $\phi_{t}$ at time $t$. As shown in Figure 2, we utilize a time-embedded UNet, frequently used in diffusion models [15, 40] to learn the mapping function. The time variable $t$ is first projected to the latent embedding space by a sinusoidal positional encoding layer and two-layer MLPs. Next, the time latent vector is fed to two separate MLPs to steer the mean and variance of the inter-layer adaptive instance normalization. The network is thus instilled with the capability of generating meaningful time embedding and predicting time-continuous flow fields by the novel time conditioning mechanism. Moreover, to ensure the predicted deformation trajectory passes through each visit $I_{k}$ at time $t=t_{k}$, we incorporate the intermediate time points as supervision in training.
3.2 Interpolation Constraints
Intermediate Similarity.
Our goal is to learn a time-continuous registration function $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$. When $t=0$, we have $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$, the identity mapping. When $t\in(0,1]$, $\phi_{t}^{0\rightarrow L}$ is the forward deformation mapping from $I_{0}$ towards $I_{L}$, and $\phi_{1}^{0\rightarrow L}$ moves $I_{0}$ to reach $I_{L}$. Similarly, for negative values, $t<0$, $\phi_{t}$ maps backward from $I_{L}$ to $I_{0}$. Exploiting bidirectional time-continuous deformations, we can formulate a symmetric similarity measure:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{t_{k}}^{0\rightarrow L},I_{k})+\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L},I_{k}),$ (1)
which deforms both images to remove the bias introduced by single-side warping [1, 30, 16]. As shown in the yellow part in Fig. 2(b), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ describe the deformation of both $I_{0}$ and $I_{L}$ to the time $t=t_{k}$. Given an acquired MRI $I_{k}$ at time $t=t_{k}$, we can simply use it to compute the symmetric intermediate similarity and thus force the network to learn plausible deformation trajectories from $I_{0}$ to $I_{L}$ that pass through intermediate images.
Generalized Intermediate Similarity.
However, in real-world clinical scenarios, many subjects will only have MRI from two time points available. In such a case, we rewrite the symmetric similarity as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L},I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}),$ (2)
where $\hat{t}$ is sampled from a uniform distribution $\mathcal{U}(0,1)$. The registration is guided properly through this synthetic variant of the symmetric similarity, even without any intermediate observation as supervision. Moreover, taking advantage of the freedom to choose $\hat{t}$, the network can learn densely on the time axis and is no longer constrained to intermediate visits at specific times in the dataset.
Extrapolation Interpolation
[…]
Seq-VXM - 4.5$\pm$0.9 25.2$\pm$1.7 2.6$\pm$0.3 0.0$\pm$0.1
(Upper bound)
rigid 10.3$\pm$1.9 16.9$\pm$1.5 - - 7.3$\pm$1.1 20.6$\pm$1.5 - -
NODER 9.3$\pm$1.9 18.0$\pm$1.7 1.8$\pm$0.9 0.2$\pm$1.9 5.5$\pm$0.9 23.2$\pm$1.5 1.8$\pm$0.5 0.0$\pm$0.3
VXM 9.1$\pm$1.9 18.3$\pm$1.9 6.3$\pm$4.8 $10^{4}\pm 10^{4}$ 5.6$\pm$1.1 22.9$\pm$1.7 1.5$\pm$0.2 0.2$\pm$0.4
VXM-diff 9.2$\pm$2.0 18.4$\pm$1.9 6.3$\pm$2.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.2$\pm$1.7 2.2$\pm$0.3 0.1$\pm$0.2
TimeFlow 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
TimeFlow-diff 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
[…]
Interpolation Consistency To encourage plausible time-continuous deformations, we propose the interpolation consistency as a regularizer:
$\begin{split}\mathcal{L}_{\textrm{inter-consis-flow}}=&\|\phi_{1}^{0\rightarrow L}\circ\phi_{t_{k}-1}^{0\rightarrow L}-\phi_{t_{k}}^{0\rightarrow L}\|_{2}^{2}\\ +&\|\phi_{-1}^{0\rightarrow L}\circ\phi_{t_{k}}^{0\rightarrow L}-\phi_{t_{k}-1}^{0\rightarrow L}\|_{2}^{2},\end{split}$ (3)
which combines the two constraints shown in Fig. 2(b). We formulated the loss with the notation $t_{k}$, but as described in the previous paragraph, $t_{k}$ can also be sampled from the uniform distribution if no intermediate observation exists. It is derived from the common transformation composition consistency that two consecutive transforms are equivalent to a direct transform. Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Fig. 2(b)). The backward case is derived similarily. It is worth noting that this differs from the commonly used smoothness regularizations, which explicitly enforce a smooth gradient over the flow field. Therefore, there is only a minimal trade-off between accuracy and smoothness since both the similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
3.3 Extrapolation Consistency
Through the intermediate similarity and interpolation consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$. As our main objective is extrapolation, we formulate criteria for the intervals $[1,+\infty)$ (forward) and $(-\infty,-1]$ (backward) next. To this end, we consider the deformation $\phi_{1}^{0\rightarrow k}$ from $I_{0}$ to $I_{k}$. Given the time of $I_{k}$ being $t_{k}$, it holds that $\phi_{t_{k}}^{0\rightarrow L}=\phi_{1}^{0\rightarrow k}$. As illustratedin Fig. 2(b), they both register $I_{0}$ to $I_{k}$. Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Fig. 2(b)), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and therefore $11/t_{k}>$ 1. Based on this notation, we define the extrapolation similarity as:
$\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{1/t_{k}}^{0\rightarrow k},I_{L}).$ (4)
Again, we can either use intermediate observations with time $t_{k}$ or sample from $\mathcal{U}(0,1)$. If $t_{k}$ is sampled from $\mathcal{U}(0,1)$, i.e., $I_{k}$ is not observed but only $I_{0}$ and $I_{L}$, we then define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$. In addition to the extrapolation similarity measure, we define the forward extrapolation flow consistency as:
$\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{1}^{0\rightarrow L}-\phi_{1/t_{k}}^{0\rightarrow k}\|_{2}^{2}.$ (5)
As shown in the bottom green part of Fig. 2(b), both $\phi_{1}^{0\rightarrow L}$ and $\phi_{1/t_{k}}^{0\rightarrow k}$ map from $I_{0}$ to $I_{L}$, thus should be consistent with each other. The backward versions of the extrapolation similarity and extrapolation flow consistency are:
$\begin{split}&\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{1/(t_{k}-1)}^{k\rightarrow L},I_{0})\quad\text{and}\\ &\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{-1}^{0\rightarrow L}-\phi_{1/(t_{k}-1)}^{k\rightarrow L}\|_{2}^{2}.\end{split}$ (6)
The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation. The proposed losses effectively ensure accurate extrapolation while implicitly capturing temporal consistency and continuity, without the need for additional temporal smoothness constraints.
4 Data and Experiments
Figure 3: Visualization of Interpolation and Extrapolation Results. TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8. The first row displays the ground truth images. The second row shows the moved baseline image. The third and fourth rows display error maps: the initial error map before registration and the error map after TimeFlow/diff registration. Error values range from $[-1,1]$ due to the image intensity normalization to $[0,1]$. These results demonstrates that TimeFlow/diff achieves high accuracy in both temporal interpolation and extrapolation. Detailed temporal deformations can be referred to Appendix.
Data and Pre-processing.
We evaluate our method on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset, which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or MCI, i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain. We employ the longitudinal processing from FreeSurfer v7.2 [13]. First, all scans are processed individually, including skull stripping, normalization. Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [39]. FreeSurfer aggregates information across time points and refines the pre-processed outputs. We first select 134 subjects who underwent substantial anatomical change (median change magnitude within the brain is $>0.3$mm) over time. Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits. It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12. All input scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels. The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping.
Implementation Details.
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $1$. We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
Baselines.
To evaluate the registration performance we compare TimeFlow with the following methods:
  • NODER [51, 3], which is aware of every time point and takes the complete image sequence as the model input. It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages Neural ordinary differential equation (ODE) [8] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [4] and the extended diffeomorphic VoxelMorph (VXMdiff[10]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The integration steps of VXMdiff is also set to 7.
  • Sequence-based VoxelMorph (Seq-VXM) [50]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point. Given that Seq-VXM performs multiple pairwise registrations by taking the whole image sequence as model input, it is aware of all images at every time point. Therefore, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab. 1).
Figure 4: Visualization of the Extrapolation Results. Left: Error maps show differences between the extrapolated predictions and the ground truth future images. Right: Future image predictions for different $t$ values, where no ground truth is available. $t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$. Thus, $t=2,4,6$ represents two, four, six-step extrapolation over $I_{f}$. Major errors are highlighted with red arrows. VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
For all learning-based methods, we apply Local Normalized Cross Correlation (LNCC) as the training loss, the same as in TimeFlow/diff. The radius of LNCC in NODER is set to $r=9$ while for the other methods, it is set to $r=4$. The loss function weighting factors in NODER follows their original suggestion. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [45, 46] with weight $1.5,0.5,1.0$.
Evaluation metrics.
[…]
Figure 5: Left: Progression Analysis for Control (CN), MCI and Dementia. A progression rate closer to 1 indicates slower pathological brain structural change. Right: Relative Progression of 9 representative subjects. Each colored dot represents one visit. The $x$-axis denotes the years since the first visit. The $y$-axis stands for the relative progression rate compared to the baseline visit. A higher value indicates faster brain structural changes. The grey line in the background is the mean fitted linear line across subjects. Colormap: CN-green, MCI-orange, Dementia-red.
5 Results
5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective studies, i.e., for extrapolating the deformations to predict follow-up images. Our model can perform this prospective analysis in two ways: 1. Based on two available scans and given a future timepoint $t$ ($t>1$), our model can directly predict the extrapolated flow field and the warped image at $t$. 2. Given a future scan, our model can estimate the corresponding $\phi_{t}$, thus the time $t$, which can reflect the subjects’ biological aging progression.
Given t, predict the future image.
As discussed in Sec. 3, one of TimeFlow’s biggest benefits is its ability to naturally extrapolate the deformation and predict a future image given $t>1$. The quantitative evaluation of extrapolation is showcased in table Tab. 1. TimeFlow demonstrates superior extrapolation accuracy compared to all baselines with the lowest MAE and the highest PSNR. Seq-VXM is not compared here since it cannot carry out future predictions. Fig. 3 (Year 8) and Fig. 4 (Left) illustrate TimeFlow’s extrapolation prediction when ground truth images are available, while Fig. 4 (Right) demonstrates the extrapolation given a large $t$ where ground truth is not available anymore. As shown in Fig. 4, NODER extrapolates unrealistic deformations, leading to significant error in future image predictions. VXM/diff, which applies linear scaling to flow fields, results in unsmooth deformation and dissolves the brain, causing anatomical details to fade. In contrast, TimeFlow/diff generates smooth and accurate extrapolated flow fields, aligning with the lowest MAE and low NDV in Tab. 1.
Given future scan, assess the brain aging.
We can utilize the extrapolation capability of TimeFlow for disease progression analysis. For each future scan, we predict a corresponding $t$ value based on prior visits. We compare the relative brain aging rate across diagnosis groups in the ADNI dataset, i.e., Controls, MCI, and Dementia. Specifically, MRI triplets are sampled from the validation and test set, where the intervals between the first and second timepoints are similar to the intervals between the second and third timepoints. The first two time points are used as input to extrapolate to the third, future time point. By comparing with the ground truth MRI, we compute the optimal extrapolated time $t_{\textrm{ext}}$. From this, we can compute the brain aging rate by dividing the extrapolated time by the actual time interval. The idea is that if brain aging is accelerated, i.e., in MCI or Dementia patients, the progression rate will be larger than 1, meaning that brain aging is faster than chronological aging. As shown in the boxplot of Fig. 5, Dementia exhibits the highest progression rate, followed by MCI, while the control group remains is close to $1$, highlighting a biologically plausible separation of groups based with our brain aging rate metric. To visualize the brain aging rate over time, we plot subject-specific progression curves for 9 representative subjects (3 from each group). We observe accelerated progression compareed to the baseline visit in MCI and Dementia, while the control group (CN) maintains a flatter curve. This highlights how TimeFlow’s extrapolation capabilities enables meaningful prospective analysis on Alzheimer’s disease, without the need for additional anatomical annotations.
$\mathcal{L}_{\textrm{inter}}$ $\mathcal{L}_{\textrm{ext}}$ Extrapolation Interpolation
[…]
1.58 $\pm$ 0.17 16.7$\pm$7.2 13.9$\pm$4.1 $>10^{3}$ $>10^{7}$ 5.3$\pm$1.0 23.3$\pm$1.7 4.7$\pm$1.3 $>10^{4}$
0.82 $\pm$ 0.18 9.8$\pm$2.2 17.6$\pm$2.1 9.7$\pm$4.8 $>10^{4}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.8$\pm$0.1 0.3$\pm$0.7
0.62 $\pm$ 0.12 8.7$\pm$1.7 18.7$\pm$1.8 5.3$\pm$1.4 $>10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
0.60 $\pm$ 0.11 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
3.25 $\pm$ 0.25 27.5$\pm$16.2 11.4$\pm$6.2 $>10^{2}$ $>10^{6}$ 5.5$\pm$1.1 22.8$\pm$1.9 6.9$\pm$2.2 $>10^{2}$
0.82 $\pm$ 0.12 10.4$\pm$2.9 17.2$\pm$2.7 8.4$\pm$2.7 $>10^{3}$ 5.3$\pm$1.0 23.4$\pm$1.8 2.5$\pm$0.2 0.3$\pm$0.7
0.55 $\pm$ 0.09 8.8$\pm$1.8 18.6$\pm$1.9 6.1$\pm$1.1 18.8$\pm$39.4 5.3$\pm$1.0 23.3$\pm$1.8 2.3$\pm$0.2 0.0$\pm$0.1
0.56 $\pm$ 0.09 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
[…]
5.2 Retrospective studies: Matching the deformation trajectory
We conduct studies on timepoints interpolation to assess whether TimeFlow learns temporally continuous and plausible deformation field connecting intermediate visits. In this setting, the baseline visit $I_{0}$ and the last scan $I_{L}$ serve as inputs, and the deformation fields $\phi_{t_{k}}$ are estimated to register intermediate images $I_{k}$, where $0<K<L$. Quantitative comparisons with baseline methods are presented in Tab. 1. As discussed in Sec. 4, Seq-VXM process the entire sequence and performs explicit pairwise registration for each timepoint, making it the performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER, which is also aware of the entire sequence and parametrized by NeuralODE. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without any extra spatial smoothness regularization. The interpolation results shown in Fig. 3 (Year 1.5) futher demonstrate that TimeFlow achieves almost zero residual errors compared to the initial misalignment.
5.3 Ablation Study
As described in Sec. 3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit temporal or spatial smoothness regularization. To assess the effectiveness of our proposed consistency losses, we conduct ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain and then take the mean over all brain voxels as the final metric. As shown in Tab. 2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically decrease) while maintaining similar interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to predict plausible extrapolated deformation fields for future image estimation. These results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields without explicit regularization, thus leading to the least trade-off between accuracy and smoothness. Visualization of the temporal deformations can be referred to Appendix.
6 Discussion and Conclusion
Novelty.
TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without temporal smoothness constraints. Together with the time-conditioning mechanism inspired by diffusion models, this enables effective extrapolation beyond the observed time frame. While we have used UNet as the base model, our contributions are generic and can, therefore, also be integrated into other registration backbones. Importantly, TimeFlow has significant application potential in areas such as in-silico drug effectiveness analysis. By extrapolating brain images to future timepoints at the beginning of a treatment, the predicted outcomes can be compared with actual follow-up scans. Discrepancies between predicted and observed results may then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
Limitation and failure cases.
TimeFlow encounters challenges in the following two scenarios: 1. Future image prediction/aging progression analysis with minimal biological aging differences: When the two base input images are temporally close and exhibit no noticeable aging differences, TimeFlow struggles to predict future images effectively based on them. In such cases, the two input images collapse into a single timepoint, rendering the model incapable of extrapolation based on a “single” input data. 2. Prediction for large time intervals: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values e.g. $t$ > 6. This limitation is expected due to the insufficient availability of long-range temporal data in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training, and, of course, the fact that modeling such long developments, which are influenced by a plethora of parameters, is inherently very challenging.
Conclusion.
In this work, we introduced TimeFlow, a novel framework for longitudinal brain MR registration that addresses key challenges in the field, including the trade-off between accuracy and temporal consistency, limited sequential data, and the inability to predict future brain states. TimeFlow leverages a UNet-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness constraints. Unlike existing methods that require full sequential data, TimeFlow learns directly from even just two images per patient, achieving both high registration accuracy and robust extrapolation to future timepoints. Our results show that TimeFlow outperforms baselines in both interpolation and extrapolation tasks and facilitates meaningful brain aging and neurodegenerative disease progression analysis. By bridging interpolation and future state prediction, TimeFlow facilitates clinically relevant analyses without extra anatomical annotations, representing a significant step toward data-driven models for tracking, predicting, and ultimately intervening in neurodegenerative diseases.
Acknowledgements
This work was supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWi) through the project NeuroTEMP, and by the Munich Center of Machine Learning (MCML). Jiazhen Pan is funded by the European Research Council (ERC) grant Deep4MI (No. 884622).
The authors declare that they have no competing interests.
Supplementary Material
7 Pairwise Registration
Pairwise Baseline Endpoint
[…]
rigid 11.0$\pm$4.5 16.6$\pm$2.2 - -
Seq-VXM 7.0$\pm$4.1 21.2$\pm$2.9 3.4$\pm$0.8 0.8$\pm$2.7
NODER 8.0$\pm$3.9 20.0$\pm$2.6 3.2$\pm$1.3 18.8$\pm$89.6
VXM 7.3$\pm$4.5 20.9$\pm$2.9 3.0$\pm$0.6 15.2$\pm$73.2
VXMdiff 6.7$\pm$3.7 21.6$\pm$2.9 4.3$\pm$1.3 33.9$\pm$175.8
TimeFlow 7.3$\pm$4.5 20.8$\pm$2.9 3.2$\pm$0.2 2.0$\pm$4.1
TimeFlowdiff 6.6$\pm$4.2 21.7$\pm$3.0 4.5$\pm$0.5 0.3$\pm$0.5
Table 3: Quantitative registration results for Pairwise Registration between baseline and endpoint. Accuracy (foreground): MAE, PSNR; Smoothness: SDlogJ, NDV.
8 Temporal Continuous Deformation Fields
Figure 6: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 7.
Figure 7: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 6.
Figure 8: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 9.
Figure 9: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 8.
Figure 10: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 11.
Figure 11: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 10.

### v3

@@
-TimeFlow: Longitudinal Brain Image Registration and Aging Progression Analysis
+IEEEexample:BSTcontrol
 Abstract
-Predicting future brain states is crucial for understanding healthy aging and neurodegenerative diseases.
-Longitudinal brain MRI registration, a cornerstone for such analyses, has long been limited by its inability to forecast future developments, reliance on extensive dense longitudinal data, and the need to balance registration accuracy with temporal smoothness.
-In this work, we present TimeFlow, a novel framework for longitudinal brain MRI registration that overcomes all these challenges.
-TimeFlow leverages a U-Net architecture with temporal conditioning inspired by diffusion models, enabling accurate registration using only two images as input and facilitating prospective analyses through future image prediction.
-Unlike traditional methods, TimeFlow eliminates the demand for explicit smoothness regularizers and dense sequential data while maintaining temporal consistency and continuity.
-Experimental results highlight its superior performance in both future timepoint prediction and registration accuracy compared to state-of-the-art methods.
-Additionally, TimeFlow supports novel biological brain aging analyses, effectively differentiating neurodegenerative conditions from healthy aging, all without requiring segmentation thus avoiding non-trivial annotation and inconsistent segmentation flaws.
-This framework paves the way for accurate, data-efficient, and annotation-free prospective analyses of brain aging and chronic diseases.
-Figure 1: Outstanding performance of TimeFlow for future brain MRI prediction.
-TimeFlow demonstrates the superior capability of future extrapolation compared to (i) NODER, which employs NeuralODE for geodesic regression, and (ii) VoxelMorph with linear scaling.
-Prediction errors in these methods are marked with red arrows.
-Given two previous timepoints (Baseline and Year 3), TimeFlow can extrapolate precise and smooth deformation to predict brain MRI in Year 6.
+Longitudinal brain analysis is essential for understanding healthy aging and identifying pathological deviations.
+Longitudinal registration of sequential brain MRI underpins such analyses.
+However, existing methods are limited by reliance on densely sampled time series, a trade-off between accuracy and temporal smoothness, and an inability to prospectively forecast future brain states.
+To overcome these challenges, we introduce TimeFlow, a learning-based framework for longitudinal brain MRI registration.
+TimeFlow uses a U-Net backbone with temporal conditioning to model neuroanatomy as a continuous function of age.
+Given only two scans from an individual, TimeFlow estimates accurate and temporally coherent deformation fields, enabling non-linear extrapolation to predict future brain states.
+This is achieved by our proposed inter-/extrapolation consistency constraints applied to both the deformation fields and deformed images.
+Remarkably, these constraints preserve temporal consistency and continuity without requiring explicit smoothness regularizers or densely sampled sequential data.
+Extensive experiments demonstrate that TimeFlow outperforms state-of-the-art methods in terms of both future timepoint forecasting and registration accuracy.
+Moreover, TimeFlow supports novel biological brain aging analyses by differentiating neurodegenerative trajectories from normal aging without requiring segmentation, thereby eliminating the need for labor-intensive annotations and mitigating segmentation inconsistency.
+TimeFlow offers an accurate, data-efficient, and annotation-free framework for longitudinal analysis of brain aging and chronic diseases, capable of forecasting brain changes beyond the observed study period.
+{IEEEkeywords}
+Brain MRI Image Registration, Longitudinal Analysis, Brain Aging Analysis
 1 Introduction
-Understanding how the human brain changes over time — both in the course of healthy aging and under the influence of neurodegenerative diseases — remains a central challenge in neurology and related fields [33, 34].
-Such insights not only shed light on the underlying pathological processes but also pave the way for more individualized treatment planning and earlier, more targeted interventions [6].
-Magnetic Resonance (MR) imaging has become the standard modality for capturing these longitudinal changes, owing to its high spatial resolution and excellent soft-tissue contrast.
-By repeatedly scanning the same individual, clinicians and researchers can visualize evolving patterns of atrophy, lesion progression, and other structural transformations as they unfold [21, 2].
-Despite advances in longitudinal registration [17, 43], which aligns a subject’s brain images across multiple time points, two critical challenges persist: 1.
-Predicting future brain states and analyzing aging progression is one of the ultimate goals of longitudinal registration.
-While retrospective assessments of past changes are valuable, the ability to predict future brain transformations would dramatically enhance clinical utility [27] and enable clinicians to forecast atrophy trajectories and guide more effective interventions [23].
-By extrapolating physically plausible transformations, longitudinal registration enables future scan prediction through simple image warping, avoiding the need for complex generative approaches.
-Moreover, comparing the predicted future brain changes with the actual observed brain state allows clinicians to assess whether the disease is advancing at an accelerated rate compared to historical progression.
-This insight can provide early warnings of disease exacerbation and help refine patient-specific prognostic models.
-Recent works [50, 3, 51] primarily focus on modeling the observed sequence of scans, allowing interpolation between observed timepoints but lacking the capability to extrapolate beyond the final scan, as illustrated in Fig.
-1.
-This limitation makes it difficult to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly.
+Understanding how the human brain evolves throughout the lifespan, whether during healthy aging or in the presence of neuropathological disorders such as neurodegenerative diseases, remains a central challenge in neurology and neuroradiology [1, 2].
+Distinguishing these trajectories not only clarifies pathological mechanisms but also refines prognostic models, ultimately supporting timely and personalized interventions [3].
+Magnetic Resonance Imaging (MRI) has become the standard tool for capturing longitudinal brain changes, owing to its high spatial resolution and superior soft-tissue contrast.
+Repeated imaging sessions allow clinicians and researchers to investigate evolving patterns of atrophy, lesion progression, and other structural alterations over time [4, 5].
+Figure 1: Top: Given two longitudinal scans (visits at Baseline and Year 4 visits), TimeFlow can interpolate and extrapolate novel brain MRI scans at unseen timepoints (e.g., Year 2 and Year 7).
+Bottom: Visual comparison of extrapolated predictions at Year 7 demonstrates TimeFlow’s superior performance, compared to (i) NODER, which uses NeuralODE-based geodesic regression, and (ii) VoxelMorph, which linearly scales displacement field.
+Prediction errors are marked by red arrows.
+At the core of most quantitative longitudinal brain analyses lies longitudinal registration, which aligns a subject’s brain scans across multiple timepoints [6, 7].
+Despite steady advances in longitudinal registration, two critical challenges persist: 1.
+Ensuring temporal continuity with limited sequential data.
+Existing methods often rely on explicit temporal smoothness constraints [8, 9, 10] which usually compromise accuracy.
+Besides, these methods often require densely sampled longitudinal data (often more than five scans) [11, 12], which is impractical in typical clinical scenarios where usually fewer than three follow-up scans are available.
 2.
-Ensuring temporal continuity with limited sequential data.
-Existing methods impose temporal smoothness constraints [49, 53, 9] and often require densely sampled longitudinal data (often more than five scans) [28, 36], which is impractical in typical clinical scenarios where only a few follow-up scans (<3) are available.
-We introduce TimeFlow, a novel longitudinal brain MRI registration capable of predicting future brain changes.
-The key contribution is a novel constraint that ensures temporal consistency and continuity — removing the demand for smoothness regularizers.
-Moreover, TimeFlow can operate longitudinal registration with only paired images in inference, eliminating the requirement for long image sequences.
-Finally, TimeFlow can extrapolate to future time points and analyze aging progression without relying on segmentation masks.
-Our contributions are twofold:
+Predicting future brain states and analyzing aging progression.
+While retrospective assessments of past changes are informative, prospective forecasting of brain anatomy would markedly enhance clinical utility [13].
+Such forecasting capabilities enable clinicians to anticipate accelerated disease progression and adjust interventions proactively [14].
+However, most recent methods model only the observed scan sequence [15, 16, 17], allowing interpolation between existing timepoints.
+It limits their capability in extrapolating beyond existing scans, as illustrated in Figure˜1, thereby restricting their ability to assess deviations from expected progression, anticipate disease worsening, and adjust treatment accordingly.
+To overcome these limitations, we introduce TimeFlow, a novel learning-based longitudinal brain MRI registration framework.
+TimeFlow combines a U-Net backbone with temporal conditioning to model neuroanatomy as a continuous function of age.
+As shown in Figure˜2, TimeFlow requires only two scans per subject to estimate the bidirectional deformation fields for arbitrary timepoints, facilitating both interpolation (timepoints within observed intervals) and extrapolation (timepoints beyond observed intervals).
+Our proposed bidirectional inter-/extra-polation consistency constraints allow unlimited sampling of timepoints, rather than the limited discrete visits during training, ensuring temporally coherent deformation fields.
+This synergy between the temporal conditioning and the proposed constraints enables TimeFlow to predict future brain states and supports aging analyses without additional annotations.
+Unlike generative models that create synthetic MRI scans, TimeFlow uses image warping to ensure anatomically plausible outputs, maintain consistency, and provide interpretability through deformation field inspection.
+Our contributions are threefold:
   1. 1.
-    Inspired by conditioning mechanisms used in diffusion models, TimeFlow employs a U-Net architecture conditioned on a temporal parameter reflecting biological aging.
-This temporal conditioning allows for generating temporally continuous and smooth deformation fields, even when only two images are used as input.
-Moreover, this approach enables future brain images prediction, surpassing state-of-the-art methods in accuracy and smoothness.
+    Temporally-conditioned registration.
+A unified network conditioned on a continuous time parameter produces temporally coherent deformation fields from only two input scans, accurately predicting unobserved anatomical states.
   2. 2.
-    We develop a novel methodology for longitudinal brain aging and disease progression analysis, utilizing TimeFlow’s advanced predictive capabilities.
-Our approach quantifies relative aging progression using healthy subjects as a reference baseline, where biological aging is assumed to align with chronological aging.
-This method enables the identification and differentiation of subjects with Mild Cognitive Impairment (MCI) and Dementia, who typically exhibit accelerated biological aging compared to chronological aging.
-Remarkably, our method performs this analysis solely through registration, eliminating the need for brain structure annotations and avoiding potential segmentation-related error propagation.
--sho
+    Bidirectional inter-/extra-polation consistency loss.
+Symmetric supervision on bidirectional deformations removes the need for explicit temporal smoothness regularizers, significantly improving temporal coherence, continuity, and generalization through unlimitedly sampled timepoints.
+  3. 3.
+    Registration-only aging analysis.
+A registration-only pipeline quantifies relative aging trajectories by referencing healthy controls, differentiating Mild Cognitive Impairment (MCI) and Dementia without requiring tissue labels, thus removing the need for non-trivial annotations.
+Figure 2: TimeFlow takes source and target images as input, conditioned on the continuous temporal parameter $t$.
+Adjusting $t$ enables the generation of interpolated brain scans within the observed interval ($0<t<1$), or extrapolated scans for future ($t>1$).
 2 Related Work
-Cross-sectional registration
-methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [41, 29], optical flow [35, 52], and diffeomorphic transformations [44, 47, 5].
+2.1 Registration across subjects
+Cross-sectional registration methods laid the groundwork for medical image registration, pioneering this field with techniques such as parametric spline-based models [18, 19], optical flow [20, 21], and diffeomorphic transformations [22, 23, 24].
 However, they demand prohibitive computing time in the order of several hours to register one single subject, posing a significant barrier to clinical deployment.
@@
 Additional tuning or adaptive strategies are often needed to strike an optimal balance [31, 32, 33].
-Longitudinal registration
-methods [42, 19, 20, 11, 12, 36, 9, 24] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject.
+2.2 Registration across time
+Longitudinal registration methods [34, 35, 36, 37, 38, 12, 10, 39] often leverage the backbone architecture of the cross-sectional registration, whereas one additional constraint is introduced: the temporal continuity/consistency across multiple timepoints of a subject.
 Two major strategies are often used to tackle this problem.
@@
 In real-world clinical scenarios, patients often have only a few follow-up scans ($\leq 3$), restricting model flexibility and practical implementation.
-Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies such as future image extrapolation and aging/disease progression risk prediction.
+Importantly, these methods commonly can only carry out retrospective analyses and are not capable of prospective studies, such as future image extrapolation and aging/disease progression risk prediction.
+2.3 Longitudinal brain analysis
+Existing longitudinal studies of neurodegenerative diseases like Alzheimer’s disease usually require additional annotations, such as cortical surfaces and structural segmentations to quantify the brain changes.
+However, these labels are usually created by cross-sectional pipelines that process each visit independently, which are both time-consuming and prone to temporal inconsistency, i.e., the same anatomy receives different labels at successive visits.
+Consequently, quantities derived from those labels (e.g., hippocampal volume, cortical-thickness trajectories) require post-processing to suppress the systematic noise [40, 41].
+These limitations motivate annotation-free, temporally coherent models that treat the sequential series as a single spatiotemporal object instead of a collection of disconnected timepoints.
+Figure 3: Left: Architecture of TimeFlow.
+The network $f_{\theta}$ incorporates temporal conditioning via adaptive instance normalization.
+It takes two input images ($I_{0}$$I_{L}$) and a continuous time variable $t$ to predict the temporal deformation field $\phi_{t}^{0\rightarrow L}$.
+Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown.
+Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency Constraints.
+The symmetric intermediate image similarity enforces the deformed source and target to be aligned at time $t_{k}$, while the interpolation flow consistency loss encourages the deformation composition consistency.
+The extrapolation similarity and flow constraints give information on what a realistic and reasonable forecast should be.
+Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$.
+The input images are highlighted by blue squares (Input), while the supervision is highlighted by dashed squares (Sup.).
+2.4 Future prediction via generative models
+Recent advances in diffusion-based generative models have inspired several attempts to synthesize longitudinal brain MRIs conditioned on variables such as age and diagnosis [42, 43].
+Despite producing visually plausible synthetic images, these methods largely remain black-box, as their diffusion trajectory does not map clearly onto biophysical deformations.
+This limits the interpretability, particularly for clinicians interested in tracking local changes like region-specific atrophy.
+Besides, diffusion models typically require large-scale datasets for stable training, which is impractical in clinical settings where follow-up is sparse.
+Nonetheless, generative models retain promise as complementary tools in neuropathological disorders involving intensity anomalies, such as gliomas or lesions, where longitudinal registration alone is insufficient to represent newly emerging tissue structures.
 3 Methods
-Figure 2: Left: Architecture of TimeFlow.
-TimeFlow $f_{\theta}$ integrates time $t$ information via adaptive instance normalization.
-The network takes the time $t$, the images at first $I_{0}$ and last visit $I_{L}$ as input, and predicts the temporal deformations.
-Examples of $t=0.5$ (interpolation at middle), $t=1$ (endpoint registration), and $t=2$ (extrapolation) are shown.
-Right: Interpolation (upper yellow) and Extrapolation (bottom green) Consistency.
-TimeFlow can predict bidirectional time-continuous deformation fields.
-The symmetric midpoint similarity forces the deformed images to be aligned at time $t_{k}$, while the midpoint interpolation flow loss encourages the deformation composition consistency.
-The extrapolation similarity and flow enforcement give information on what a realistic and reasonable extrapolation should be.
-Training with both consistency losses enables TimeFlow to generate plausible deformation in $t\in[-\infty,+\infty]$.
-The input images are highlighted by blue squares (Input), while the supervision are highlighted by dot-lined squares (Sup.).
 3.1 Time-Continuous Registration
-Given a sequence of longitudinal images $\{I_{k}\mid k=0,\ldots,L\}$, we aim to find a temporally continuous deformation trajectory that connects all the discrete time points.
-To achieve this, we construct a time-continuous registration function $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$.
-It takes the first visit $I_{0}$ and the last visit $I_{L}$ as input images and predicts the deformation $\phi_{t}$ at time $t$.
-As shown in Figure 2, we utilize a time-embedded UNet, frequently used in diffusion models [15, 40] to learn the mapping function.
-The time variable $t$ is first projected to the latent embedding space by a sinusoidal positional encoding layer and two-layer MLPs.
+Given a longitudinal MRI sequence $\{I_{k}\mid k=0,\ldots,L\}$, our goal is to derive a temporally continuous deformation trajectory connecting all intermediate timepoints.
+To achieve this, we propose a time-continuous registration framework, TimeFlow, defined as $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$.
+As illustrated in Figure˜2, TimeFlow accepts the baseline image $I_{0}$ (source) and the last image $I_{L}$ (target) and predicts the deformation fields $\phi_{t}$ parametrized by continuous time $t$.
+For interpolation within the observed interval ($0<t<1$), TimeFlow predicts intermediate states between source and target scans.
+For extrapolation ($t>1$), TimeFlow predicts future brain states beyond the last observed visit.
+To enforce temporal coherence and continuity in the deformation trajectories, we introduce the novel inter-/extra-polation consistency constraints.
+The core architecture (Figure˜3, Left) is a time-embedded U-Net [44, 45].
+Specifically, a sinusoidal positional encoding and multi-layer perceptrons (MLPs) project the scalar time variable $t$ into a latent embedding space.
 Next, the time latent vector is fed to two separate MLPs to steer the mean and variance of the inter-layer adaptive instance normalization.
 The network is thus instilled with the capability of generating meaningful time embedding and predicting time-continuous flow fields by the novel time conditioning mechanism.
-Moreover, to ensure the predicted deformation trajectory passes through each visit $I_{k}$ at time $t=t_{k}$, we incorporate the intermediate time points as supervision in training.
-3.2 Interpolation Constraints
-Intermediate Similarity.
-Our goal is to learn a time-continuous registration function $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$.
-When $t=0$, we have $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$, the identity mapping.
+3.2 Interpolation Consistency
+Symmetric Intermediate Similarity
+We first define the time-continuous registration function as $\phi_{t}^{0\rightarrow L}=f_{\theta}(I_{0},I_{L},t)$, where $f_{\theta}(I_{0},I_{L},0)=\textrm{Id}$ is the identity mapping at $t=0$.
 When $t\in(0,1]$, $\phi_{t}^{0\rightarrow L}$ is the forward deformation mapping from $I_{0}$ towards $I_{L}$, and $\phi_{1}^{0\rightarrow L}$ moves $I_{0}$ to reach $I_{L}$.
 Similarly, for negative values, $t<0$, $\phi_{t}$ maps backward from $I_{L}$ to $I_{0}$.
-Exploiting bidirectional time-continuous deformations, we can formulate a symmetric similarity measure:
+For any intermediate image $I_{k}$ observed at time $t_{k}$, we can formulate the symmetric interpolation image similarity measure as:
 $\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{t_{k}}^{0\rightarrow L},I_{k})+\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L},I_{k}),$ (1)
-which deforms both images to remove the bias introduced by single-side warping [1, 30, 16].
-As shown in the yellow part in Fig.
-2(b), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ describe the deformation of both $I_{0}$ and $I_{L}$ to the time $t=t_{k}$.
-Given an acquired MRI $I_{k}$ at time $t=t_{k}$, we can simply use it to compute the symmetric intermediate similarity and thus force the network to learn plausible deformation trajectories from $I_{0}$ to $I_{L}$ that pass through intermediate images.
-Generalized Intermediate Similarity.
-However, in real-world clinical scenarios, many subjects will only have MRI from two time points available. In such a case, we rewrite the symmetric similarity as:
+which deforms both source and target images to remove the bias introduced by single-side warping [46, 47, 48].
+As shown in Figure˜3 (Right, top yellow), $I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ is warping $I_{0}$ to the time $t=t_{k}$ and $I_{L}\circ\phi_{t_{k}-1}^{0\rightarrow L}$ is warping $I_{L}$ to the time $t=t_{k}$.
+This symmetry enforces bidirectional consistency, facilitating plausible deformation trajectories passing through observed intermediate scans.
+Generalized Symmetric Intermediate Similarity
+However, in real-world clinical scenarios, many subjects have fewer than three MRI scans available.
+In such a case, we generalize the above constraint by synthetically sampling intermediate timepoints $\hat{t}$ from a uniform distribution $\mathcal{U}(0,1)$, defining the generalized interpolation similarity as:
 $\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L},I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}),$ (2)
-where $\hat{t}$ is sampled from a uniform distribution $\mathcal{U}(0,1)$.
-The registration is guided properly through this synthetic variant of the symmetric similarity, even without any intermediate observation as supervision.
-Moreover, taking advantage of the freedom to choose $\hat{t}$, the network can learn densely on the time axis and is no longer constrained to intermediate visits at specific times in the dataset.
-Extrapolation Interpolation
-[…]
-Seq-VXM - 4.5$\pm$0.9 25.2$\pm$1.7 2.6$\pm$0.3 0.0$\pm$0.1
-(Upper bound)
-rigid 10.3$\pm$1.9 16.9$\pm$1.5 - - 7.3$\pm$1.1 20.6$\pm$1.5 - -
-NODER 9.3$\pm$1.9 18.0$\pm$1.7 1.8$\pm$0.9 0.2$\pm$1.9 5.5$\pm$0.9 23.2$\pm$1.5 1.8$\pm$0.5 0.0$\pm$0.3
-VXM 9.1$\pm$1.9 18.3$\pm$1.9 6.3$\pm$4.8 $10^{4}\pm 10^{4}$ 5.6$\pm$1.1 22.9$\pm$1.7 1.5$\pm$0.2 0.2$\pm$0.4
-VXM-diff 9.2$\pm$2.0 18.4$\pm$1.9 6.3$\pm$2.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.2$\pm$1.7 2.2$\pm$0.3 0.1$\pm$0.2
-TimeFlow 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
-TimeFlow-diff 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
-[…]
-Interpolation Consistency To encourage plausible time-continuous deformations, we propose the interpolation consistency as a regularizer:
+Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timpoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations.
+By freely sampling $\hat{t}$, TimeFlow effectively learns densely and continuously across the entire temporal axis.
+Interpolation Flow Consistency
+To further encourage plausible time-continuous deformations, we propose the interpolation flow consistency as a regularizer:
 $\begin{split}\mathcal{L}_{\textrm{inter-consis-flow}}=&\|\phi_{1}^{0\rightarrow L}\circ\phi_{t_{k}-1}^{0\rightarrow L}-\phi_{t_{k}}^{0\rightarrow L}\|_{2}^{2}\\ +&\|\phi_{-1}^{0\rightarrow L}\circ\phi_{t_{k}}^{0\rightarrow L}-\phi_{t_{k}-1}^{0\rightarrow L}\|_{2}^{2},\end{split}$ (3)
-which combines the two constraints shown in Fig.
-2(b).
+which combines the two constraints shown in Figure˜3 (Right, top yellow).
 We formulated the loss with the notation $t_{k}$, but as described in the previous paragraph, $t_{k}$ can also be sampled from the uniform distribution if no intermediate observation exists.
 It is derived from the common transformation composition consistency that two consecutive transforms are equivalent to a direct transform.
-Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Fig.
-2(b)).
-The backward case is derived similarily.
+Moving from $t=0$ to $t=1$ and then back to $t=t_{k}$ is equivalent to moving directly from $t=0$ to $t=t_{k}$ (arrow 1 composed with arrow 3 should equal to arrow 2 in Figure˜3, Right).
+The backward case is derived similarly.
 It is worth noting that this differs from the commonly used smoothness regularizations, which explicitly enforce a smooth gradient over the flow field.
-Therefore, there is only a minimal trade-off between accuracy and smoothness since both the similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
+Therefore, there is only a minimal trade-off between accuracy and smoothness since both the symmetric image similarity and the flow consistency encourage plausible and reasonable longitudinal transformations.
 3.3 Extrapolation Consistency
-Through the intermediate similarity and interpolation consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$.
-As our main objective is extrapolation, we formulate criteria for the intervals $[1,+\infty)$ (forward) and $(-\infty,-1]$ (backward) next.
+Through the intermediate similarity and interpolation flow consistency, we obtain meaningful predictions for bidirectional registrations in $t\in[-1,1]$.
+Now we introduce extrapolation constraints for forward ($t>1$) and backward (t<1t<-1) directions.
 To this end, we consider the deformation $\phi_{1}^{0\rightarrow k}$ from $I_{0}$ to $I_{k}$.
 Given the time of $I_{k}$ being $t_{k}$, it holds that $\phi_{t_{k}}^{0\rightarrow L}=\phi_{1}^{0\rightarrow k}$.
-As illustratedin Fig.
-2(b), they both register $I_{0}$ to $I_{k}$.
-Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Fig.
-2(b)), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and therefore $11/t_{k}>$ 1.
-Based on this notation, we define the extrapolation similarity as:
+As illustrated in Figure˜3 (Right), they both register $I_{0}$ to $I_{k}$.
+Consequently, we can then formulate $\phi_{1}^{0\rightarrow L}=\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 1 equals to arrow 5 in Figure˜3, Right), where $\phi_{1/t_{k}}^{0\rightarrow k}$ denotes an extrapolation because $t_{k}\in(0,1)$ and $1/t_{k}>1$.
+Based on this notation, we define the extrapolation image similarity as:
 $\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{1/t_{k}}^{0\rightarrow k},I_{L}).$ (4)
-Again, we can either use intermediate observations with time $t_{k}$ or sample from $\mathcal{U}(0,1)$.
-If $t_{k}$ is sampled from $\mathcal{U}(0,1)$, i.e., $I_{k}$ is not observed but only $I_{0}$ and $I_{L}$, we then define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$.
+Again, if no intermediate observations are available, we can sample $t_{k}$ from $\mathcal{U}(0,1)$ and define $I_{k}=I_{0}\circ\phi_{t_{k}}^{0\rightarrow L}$ to create synthetic intermediate data.
 In addition to the extrapolation similarity measure, we define the forward extrapolation flow consistency as:
 $\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{1}^{0\rightarrow L}-\phi_{1/t_{k}}^{0\rightarrow k}\|_{2}^{2}.$ (5)
-As shown in the bottom green part of Fig.
-2(b), both $\phi_{1}^{0\rightarrow L}$ and $\phi_{1/t_{k}}^{0\rightarrow k}$ map from $I_{0}$ to $I_{L}$, thus should be consistent with each other.
-The backward versions of the extrapolation similarity and extrapolation flow consistency are:
+As shown in Figure˜3 (Right), both $\phi_{1}^{0\rightarrow L}$ (arrow 1) and $\phi_{1/t_{k}}^{0\rightarrow k}$ (arrow 5) map from $I_{0}$ to $I_{L}$, thus should be consistent with each other.
+The backward versions of the extrapolation similarity and extrapolation flow consistency are similarly defined as:
 $\begin{split}&\mathcal{L}_{\textrm{sim}}(I_{L}\circ\phi_{1/(t_{k}-1)}^{k\rightarrow L},I_{0})\quad\text{and}\\ &\mathcal{L}_{\textrm{ext-consis-flow}}=\|\phi_{-1}^{0\rightarrow L}-\phi_{1/(t_{k}-1)}^{k\rightarrow L}\|_{2}^{2}.\end{split}$ (6)
-The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation.
+The similarity measures and flow consistencies both contribute to achieving a realistic and coherent extrapolation forecasting.
 The proposed losses effectively ensure accurate extrapolation while implicitly capturing temporal consistency and continuity, without the need for additional temporal smoothness constraints.
+3.4 Unlimited Sampled Triplets
+As illustrated in Figure˜3 (Right), the proposed inter-/extra-polation constraints require observed triplets of timepoints, limiting learning when few discrete scans are available.
+To address this, we use “synthetic” triplets dynamically sampled from continuous time intervals.
+Specifically, by sampling intermediate timepoints $\hat{t}$ from $\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints.
+This continuous, densely sampled temporal supervision significantly enhances the network’s capacity to generalize and accurately model brain anatomy changes at any arbitrary timepoint, including unseen future scenarios, while using only two scans as inputs.
 4 Data and Experiments
-Figure 3: Visualization of Interpolation and Extrapolation Results.
-TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8.
-The first row displays the ground truth images.
-The second row shows the moved baseline image.
-The third and fourth rows display error maps: the initial error map before registration and the error map after TimeFlow/diff registration.
-Error values range from $[-1,1]$ due to the image intensity normalization to $[0,1]$.
-These results demonstrates that TimeFlow/diff achieves high accuracy in both temporal interpolation and extrapolation.
-Detailed temporal deformations can be referred to Appendix.
-Data and Pre-processing.
-We evaluate our method on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset, which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or MCI, i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain.
+Data and Pre-processing
+We train and evaluate TimeFlow on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [4], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain.
 We employ the longitudinal processing from FreeSurfer v7.2 [49].
@@
 FreeSurfer aggregates information across time points and refines the pre-processed outputs.
-We first select 134 subjects who underwent substantial anatomical change (median change magnitude within the brain is $>0.3$mm) over time.
+To ensure meaningful longitudinal analysis, we select 134 subjects exhibiting substantial structural changes (median brain deformation >0.3>0.3mm).
 Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits.
 It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12.
-All input scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels.
+All scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels.
 The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping.
-Implementation Details.
+Additionally, 11 subjects from the OASIS dataset [51] serve for zero-shot evaluation.
+Figure 4: Visualization of extrapolation errors compared to ground truth future images. Coronal and axial views of two subjects are presented. Areas with high error are marked by red arrows.
+Implementation Details
 Our model is implemented using Pytorch 2.1.2.
-and training/test is carried out on one NVIDIA A100 40G with a batch size of $1$.
+and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per subject).
 We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$.
@@
 The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
-Baselines.
-To evaluate the registration performance we compare TimeFlow with the following methods:
-  • NODER [51, 3], which is aware of every time point and takes the complete image sequence as the model input.
-It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages Neural ordinary differential equation (ODE) [8] to compute integration and adjoint gradients.
+Figure 5: Future image predictions at different extrapolation time $t$, where no ground truth is available.
+$t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$.
+Thus, $t=2,4,6$ represents two-, four-, and six-step extrapolation over $I_{f}$.
+Major errors are highlighted with red arrows.
+VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
+[…]
+Extrapolation Interpolation
+[…]
+[…]
+(Upper bound)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 6: Visualization of interpolation (Year 1.5) and extrapolation (Year 8) results from coronal and axial views of a subject’s brain.
+TimeFlow/diff takes the Baseline and Year 4 images as input, interpolates to Year 1.5, and extrapolates to Year 8.
+Ground truth images (row 1) and warped images (row 2) are compared.
+Initial misalignment (row 3) is effectively corrected (row 4), demonstrating TimeFlow’s accuracy in temporal modeling.
+Error values range from [1,1][-1,1] due to the image intensity normalization to $[0,1]$.
+Baselines
+To evaluate the registration performance, we compare TimeFlow with the following methods:
+  • NODER [17, 16] takes the complete image sequence as the model input and is thus aware of every time point.
+It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [52] to compute integration and adjoint gradients.
 It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
-  • VoxelMorph (VXM) [4] and the extended diffeomorphic VoxelMorph (VXMdiff[10].
+  • VoxelMorph (VXM) [25] and the diffeomorphic variant (VXMdiff[53].
 Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration.
-The integration steps of VXMdiff is also set to 7.
+The number of integration steps of VXMdiff is also set to 7.
   • Sequence-based VoxelMorph (Seq-VXM) [15].
 Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point.
-Given that Seq-VXM performs multiple pairwise registrations by taking the whole image sequence as model input, it is aware of all images at every time point.
-Therefore, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab.
+Given that Seq-VXM performs multiple pairwise registrations by taking the full sequence as model input, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab.
 1).
-Figure 4: Visualization of the Extrapolation Results.
-Left: Error maps show differences between the extrapolated predictions and the ground truth future images.
-Right: Future image predictions for different $t$ values, where no ground truth is available.
-$t=1$ represents one-step pairwise registration between $I_{m}$ and $I_{f}$.
-Thus, $t=2,4,6$ represents two, four, six-step extrapolation over $I_{f}$.
-Major errors are highlighted with red arrows.
-VXM/diff produces unsmooth deformations, leading to structural inconsistencies and loss of anatomical details (zoom in for more details).
-For all learning-based methods, we apply Local Normalized Cross Correlation (LNCC) as the training loss, the same as in TimeFlow/diff.
-The radius of LNCC in NODER is set to $r=9$ while for the other methods, it is set to $r=4$.
-The loss function weighting factors in NODER follows their original suggestion.
+All methods use Local Normalized Cross Correlation (LNCC) with radius $r=9$ (NODER) or $r=4$ (others) as image similarity loss.
+The loss function weighting factors in NODER follow their original suggestion.
 Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [54, 55] with weight $1.5,0.5,1.0$.
-Evaluation metrics.
+Figure 7: Left: Prospective, annotation-free analysis of brain aging progression rates for Control (CN), MCI and Dementia.
+Rates near 1 reflect normal, chronological brain aging, whereas values above 1 indicate accelerated neurodegenerative progression.
+Right: Retrospective group-wise analysis of longitudinal brain progression.
+Trajectories represent median linear fits between TimeFlow’s inferred temporal brain state ($t$) and chronological years since baseline.
+Shaded ribbons indicate the 40th-60th percentile range.
+A steeper slope corresponds to faster neuroanatomical changes.
+Colormap: CN-green, MCI-orange, Dementia-red.
+Figure 8: Quantitative comparison of extrapolation accuracy among TimeFlow and baseline methods on ADNI dataset.
+Barplots represent mean values with standard errors.
+Significance levels by Wilcoxon $t$-test (Holm–Bonferroni correction) are indicated.
+***:p<0.001, **:p<0.01, *:p<0.05
+Evaluation metrics
 […]
-Figure 5: Left: Progression Analysis for Control (CN), MCI and Dementia.
-A progression rate closer to 1 indicates slower pathological brain structural change.
-Right: Relative Progression of 9 representative subjects.
-Each colored dot represents one visit.
-The $x$-axis denotes the years since the first visit.
-The $y$-axis stands for the relative progression rate compared to the baseline visit.
-A higher value indicates faster brain structural changes.
-The grey line in the background is the mean fitted linear line across subjects.
-Colormap: CN-green, MCI-orange, Dementia-red.
 5 Results
 5.1 Prospective studies: Predicting the future
-We first evaluate TimeFlow’s performance in prospective studies, i.e., for extrapolating the deformations to predict follow-up images.
-Our model can perform this prospective analysis in two ways: 1.
-Based on two available scans and given a future timepoint $t$ ($t>1$), our model can directly predict the extrapolated flow field and the warped image at $t$.
-2.
-Given a future scan, our model can estimate the corresponding $\phi_{t}$, thus the time $t$, which can reflect the subjects’ biological aging progression.
-Given t, predict the future image.
-As discussed in Sec.
-3, one of TimeFlow’s biggest benefits is its ability to naturally extrapolate the deformation and predict a future image given $t>1$.
-The quantitative evaluation of extrapolation is showcased in table Tab.
-1.
-TimeFlow demonstrates superior extrapolation accuracy compared to all baselines with the lowest MAE and the highest PSNR.
-Seq-VXM is not compared here since it cannot carry out future predictions.
-Fig.
-3 (Year 8) and Fig.
-4 (Left) illustrate TimeFlow’s extrapolation prediction when ground truth images are available, while Fig.
-4 (Right) demonstrates the extrapolation given a large $t$ where ground truth is not available anymore.
-As shown in Fig.
-4, NODER extrapolates unrealistic deformations, leading to significant error in future image predictions.
-VXM/diff, which applies linear scaling to flow fields, results in unsmooth deformation and dissolves the brain, causing anatomical details to fade.
-In contrast, TimeFlow/diff generates smooth and accurate extrapolated flow fields, aligning with the lowest MAE and low NDV in Tab.
-1.
-Given future scan, assess the brain aging.
-We can utilize the extrapolation capability of TimeFlow for disease progression analysis.
-For each future scan, we predict a corresponding $t$ value based on prior visits.
-We compare the relative brain aging rate across diagnosis groups in the ADNI dataset, i.e., Controls, MCI, and Dementia.
-Specifically, MRI triplets are sampled from the validation and test set, where the intervals between the first and second timepoints are similar to the intervals between the second and third timepoints.
-The first two time points are used as input to extrapolate to the third, future time point.
-By comparing with the ground truth MRI, we compute the optimal extrapolated time $t_{\textrm{ext}}$.
-From this, we can compute the brain aging rate by dividing the extrapolated time by the actual time interval.
-The idea is that if brain aging is accelerated, i.e., in MCI or Dementia patients, the progression rate will be larger than 1, meaning that brain aging is faster than chronological aging.
-As shown in the boxplot of Fig.
-5, Dementia exhibits the highest progression rate, followed by MCI, while the control group remains is close to $1$, highlighting a biologically plausible separation of groups based with our brain aging rate metric.
-To visualize the brain aging rate over time, we plot subject-specific progression curves for 9 representative subjects (3 from each group).
-We observe accelerated progression compareed to the baseline visit in MCI and Dementia, while the control group (CN) maintains a flatter curve.
-This highlights how TimeFlow’s extrapolation capabilities enables meaningful prospective analysis on Alzheimer’s disease, without the need for additional anatomical annotations.
+We first evaluate TimeFlow’s performance in prospective scenarios, specifically, its ability to extrapolate deformation fields to predict future brain scans.
+TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a surrogate metric to quantify the biological aging progression.
+Given t, predict the future image
+A significant advantage of TimeFlow, as described in Section˜3, is its ability to extrapolate deformation fields to future timepoints beyond the observed interval.
+The quantitative performance of TimeFlow’s extrapolation forecasting is reported in Table˜1.
+The metrics are computed with respect to the ground truth future scans.
+Results clearly demonstrate that TimeFlow consistently achieves superior extrapolation accuracy compared to baseline methods, yielding the lowest MAE and highest PSNR (Figure˜8), even when evaluated in a zero-shot manner on the unseen OASIS dataset.
+Note that Seq-VXM cannot carry out future predictions; thus, it is excluded from this comparison.
+Qualitative extrapolation results are visualized in Figure˜6 (Year 8) and Figure˜4, illustrating TimeFlow’s accuracy by comparing to the ground truth images.
+Moreover, Figure˜5 demonstrates extrapolation at a large $t$ ($t=2,4,6$) where ground truth is not available.
+In these cases, NODER produces unrealistic and anatomically implausible deformations, resulting in significant errors in future image prediction.
+Similarly, VXM/diff, which applies linear scaling to displacement/velocity fields, generates unsmooth deformations (indicated by high NDV) and dissolves the brain, causing the anatomical details to fade.
+In contrast, TimeFlow/diff consistently generates smooth and coherent extrapolated deformation fields, quantitatively supported by the lowest MAE and minimal NDV values in Table˜1.
+Given future scan, assess the brain aging
+As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis.
+We compare the brain aging rates across diagnostic groups, i.e., Controls (CN), MCI, and Dementia, by predicting the optimal time parameter ($t_{\textrm{ext}}$) corresponding to the observed future MRI scans.
+Specifically, MRI triplets are sampled from validation and test subsets, where the intervals between the first and second visits match closely the intervals between the second and third visits.
+The diagnosis at the third visit labels each triplet.
+Using the first two scans as input, TimeFlow extrapolates forward to the third, future scan.
+By comparing predictions against actual observed future, we compute the extrapolated time parameter $t_{\textrm{ext}}$, and divide it by the normalized chronological intervals (in years) to yield a brain aging rate.
+An aging rate above 1 indicates accelerated pathological changes, as expected in patients with neurodegenerative conditions.
+Indeed, the boxplot of Figure˜7 (Left) clearly shows that Dementia patients exhibit the highest progression rate, followed by MCI subjects, while the control group maintains progression rates around 1.
+The analysis validates that our annotation-free surrogate brain aging rate metric is able to separate disease groups.
+[…]
 $\mathcal{L}_{\textrm{inter}}$ $\mathcal{L}_{\textrm{ext}}$ Extrapolation Interpolation
 […]
-1.58 $\pm$ 0.17 16.7$\pm$7.2 13.9$\pm$4.1 $>10^{3}$ $>10^{7}$ 5.3$\pm$1.0 23.3$\pm$1.7 4.7$\pm$1.3 $>10^{4}$
-0.82 $\pm$ 0.18 9.8$\pm$2.2 17.6$\pm$2.1 9.7$\pm$4.8 $>10^{4}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.8$\pm$0.1 0.3$\pm$0.7
-0.62 $\pm$ 0.12 8.7$\pm$1.7 18.7$\pm$1.8 5.3$\pm$1.4 $>10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
-0.60 $\pm$ 0.11 8.6$\pm$1.6 18.8$\pm$1.7 5.0$\pm$1.0 $10^{3}\pm 10^{3}$ 5.5$\pm$1.0 23.0$\pm$1.7 1.6$\pm$0.1 0.1$\pm$0.4
-3.25 $\pm$ 0.25 27.5$\pm$16.2 11.4$\pm$6.2 $>10^{2}$ $>10^{6}$ 5.5$\pm$1.1 22.8$\pm$1.9 6.9$\pm$2.2 $>10^{2}$
-0.82 $\pm$ 0.12 10.4$\pm$2.9 17.2$\pm$2.7 8.4$\pm$2.7 $>10^{3}$ 5.3$\pm$1.0 23.4$\pm$1.8 2.5$\pm$0.2 0.3$\pm$0.7
-0.55 $\pm$ 0.09 8.8$\pm$1.8 18.6$\pm$1.9 6.1$\pm$1.1 18.8$\pm$39.4 5.3$\pm$1.0 23.3$\pm$1.8 2.3$\pm$0.2 0.0$\pm$0.1
-0.56 $\pm$ 0.09 8.8$\pm$1.8 18.7$\pm$1.9 6.0$\pm$1.0 6.1$\pm$11.2 5.3$\pm$1.0 23.3$\pm$1.8 2.2$\pm$0.2 0.0$\pm$0.0
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 […]
 5.2 Retrospective studies: Matching the deformation trajectory
-We conduct studies on timepoints interpolation to assess whether TimeFlow learns temporally continuous and plausible deformation field connecting intermediate visits.
-In this setting, the baseline visit $I_{0}$ and the last scan $I_{L}$ serve as inputs, and the deformation fields $\phi_{t_{k}}$ are estimated to register intermediate images $I_{k}$, where $0<K<L$.
-Quantitative comparisons with baseline methods are presented in Tab.
-1.
-As discussed in Sec.
-4, Seq-VXM process the entire sequence and performs explicit pairwise registration for each timepoint, making it the performance upper bound.
-Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER, which is also aware of the entire sequence and parametrized by NeuralODE.
-Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without any extra spatial smoothness regularization.
-The interpolation results shown in Fig.
-3 (Year 1.5) futher demonstrate that TimeFlow achieves almost zero residual errors compared to the initial misalignment.
+We further evaluate TimeFlow’s interpolation capability, assessing its effectiveness in modeling temporally continuous and anatomically plausible deformation trajectories connecting observed MRI visits.
+In this setting, TimeFlow uses only the baseline visit $I_{0}$ and the final visit $I_{L}$ as inputs, predicting deformation fields $\phi_{t_{k}}$ for intermediate images $I_{k}$, where $0<k<L$.
+Quantitative comparisons with baseline methods for interpolation are presented in Table˜1.
+As discussed in baselines, Seq-VXM, performing explicit pairwise registrations for every available timepoint, represents a performance upper bound.
+Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER on the ADNI dataset, which is also aware of the entire sequence and parametrized by Neuralordinary differential equation (ODE).
+TimeFlow achieves comparable performance to NODER in zero-shot evaluation on the OASIS dataset, where NODER performs subject-specific optimization on the full image sequence.
+Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without extra spatial smoothness regularization.
+Qualitative visualizations in Figure˜6 (Year 1.5) illustrate the accuracy of TimeFlow’s interpolations, substantially reducing alignment errors compared to initial misalignments.
+Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a surrogate brain progression state marker.
+We extract trajectory segments with the same diagnoses from each subject, and then compute a linear relationship between the estimated brain state ($t$) and chronological age (years since baseline of the segment).
+The median linear lines of the groups (CN, MCI, Dementia) in Figure˜7 (Right) clearly illustrate distinct progression patterns, with Dementia exhibiting the fastest structural changes, MCI intermediate, and CN the slowest.
 5.3 Ablation Study
-As described in Sec.
-3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit temporal or spatial smoothness regularization.
-To assess the effectiveness of our proposed consistency losses, we conduct ablation studies.
+As described in Section˜3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit smoothness regularization.
+To assess the effectiveness of our proposed inter-/extra-polation consistency constraints, we perform detailed ablation studies.
 To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$.
 Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$.
-Specifically, we compute the norm of the temporal displacement derivatives within the brain and then take the mean over all brain voxels as the final metric.
-As shown in Tab.
-2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically decrease) while maintaining similar interpolation accuracy.
-Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to predict plausible extrapolated deformation fields for future image estimation.
-These results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields without explicit regularization, thus leading to the least trade-off between accuracy and smoothness.
-Visualization of the temporal deformations can be referred to Appendix.
-6 Discussion and Conclusion
-Novelty.
-TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without temporal smoothness constraints.
-Together with the time-conditioning mechanism inspired by diffusion models, this enables effective extrapolation beyond the observed time frame.
-While we have used UNet as the base model, our contributions are generic and can, therefore, also be integrated into other registration backbones.
-Importantly, TimeFlow has significant application potential in areas such as in-silico drug effectiveness analysis.
+Specifically, we compute the norm of the temporal displacement derivatives within the brain regions and then take the mean over all brain voxels as the final metric.
+As shown in table˜2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy.
+Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to generate plausible extrapolated deformation fields for future image estimation.
+These ablation results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields, achieving an optimal balance between registration accuracy and smoothness without explicit smoothness penalties.
+6 Discussion
+Contributions
+TimeFlow introduces novel interpolation and extrapolation constraints that ensure temporal consistency without explicit temporal smoothness regularization.
+Combined with the time-conditioning mechanism, TimeFlow enables effective extrapolation beyond the observed time frame.
+While we have used U-Net as the base model, our contributions are generic and can be integrated into other registration backbones.
+Importantly, TimeFlow has promising application potential in areas such as in silico drug effectiveness analysis.
 By extrapolating brain images to future timepoints at the beginning of a treatment, the predicted outcomes can be compared with actual follow-up scans.
-Discrepancies between predicted and observed results may then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
-Limitation and failure cases.
+Discrepancies between predicted and observed results could then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
+Bidirectional extrapolation
+While we mainly focus on the forward extrapolation (predicting future brain states) in this study, TimeFlow is inherently bidirectional and allows backward extrapolation as well.
+Backward extrapolation, i.e., traveling back in time and visualizing the brain at earlier, unobserved ages, opens several avenues: (i) reconstructing a pre-symptomatic baseline for patients before neurodegeneration; (ii) generating counterfactual controls that share a subject’s anatomy but simulate a healthy aging trajectory.
+Limitation and failure cases
 TimeFlow encounters challenges in the following two scenarios: 1.
-Future image prediction/aging progression analysis with minimal biological aging differences: When the two base input images are temporally close and exhibit no noticeable aging differences, TimeFlow struggles to predict future images effectively based on them.
-In such cases, the two input images collapse into a single timepoint, rendering the model incapable of extrapolation based on a “single” input data.
+Future image prediction/aging progression analysis with minimal biological aging differences: When the two input scans are temporally close and exhibit no noticeable structural differences, they collapse into a single timepoint, which TimeFlow struggles to predict future images effectively.
 2.
-Prediction for large time intervals: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values e.g.
-$t$ > 6.
-This limitation is expected due to the insufficient availability of long-range temporal data in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training, and, of course, the fact that modeling such long developments, which are influenced by a plethora of parameters, is inherently very challenging.
-Conclusion.
-In this work, we introduced TimeFlow, a novel framework for longitudinal brain MR registration that addresses key challenges in the field, including the trade-off between accuracy and temporal consistency, limited sequential data, and the inability to predict future brain states.
-TimeFlow leverages a UNet-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness constraints.
-Unlike existing methods that require full sequential data, TimeFlow learns directly from even just two images per patient, achieving both high registration accuracy and robust extrapolation to future timepoints.
-Our results show that TimeFlow outperforms baselines in both interpolation and extrapolation tasks and facilitates meaningful brain aging and neurodegenerative disease progression analysis.
-By bridging interpolation and future state prediction, TimeFlow facilitates clinically relevant analyses without extra anatomical annotations, representing a significant step toward data-driven models for tracking, predicting, and ultimately intervening in neurodegenerative diseases.
-Acknowledgements
-This work was supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWi) through the project NeuroTEMP, and by the Munich Center of Machine Learning (MCML).
-Jiazhen Pan is funded by the European Research Council (ERC) grant Deep4MI (No.
-884622).
-The authors declare that they have no competing interests.
-Supplementary Material
-7 Pairwise Registration
-Pairwise Baseline Endpoint
-[…]
-rigid 11.0$\pm$4.5 16.6$\pm$2.2 - -
-Seq-VXM 7.0$\pm$4.1 21.2$\pm$2.9 3.4$\pm$0.8 0.8$\pm$2.7
-NODER 8.0$\pm$3.9 20.0$\pm$2.6 3.2$\pm$1.3 18.8$\pm$89.6
-VXM 7.3$\pm$4.5 20.9$\pm$2.9 3.0$\pm$0.6 15.2$\pm$73.2
-VXMdiff 6.7$\pm$3.7 21.6$\pm$2.9 4.3$\pm$1.3 33.9$\pm$175.8
-TimeFlow 7.3$\pm$4.5 20.8$\pm$2.9 3.2$\pm$0.2 2.0$\pm$4.1
-TimeFlowdiff 6.6$\pm$4.2 21.7$\pm$3.0 4.5$\pm$0.5 0.3$\pm$0.5
-Table 3: Quantitative registration results for Pairwise Registration between baseline and endpoint. Accuracy (foreground): MAE, PSNR; Smoothness: SDlogJ, NDV.
-8 Temporal Continuous Deformation Fields
-Figure 6: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 7.
-Figure 7: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 6.
-Figure 8: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 9.
-Figure 9: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 8.
-Figure 10: Visualization of temporal deformations $\phi_{t}$. $t$ is sampled from $t=0.5$ to $t=2.5$ with an interval $\Delta t=0.75$. The warped moving images are shown in Fig. 11.
-Figure 11: Visualization of warped moving Images $I_{m}\circ\phi$ by the temporal deforamtons. The corresponding flow fields are shown in Fig. 10.
+Prediction for far future: Although TimeFlow is designed to estimate future images for any given time parameter $t\in[0,\infty]$, its performance deteriorates for large $t$ values, e.g., $t$ > 6.
+This limitation is expected due to limited data covering such long-range time intervals in the ADNI dataset, meaning TimeFlow rarely encountered such scenarios during training.
+The modeling of such long developments, which are influenced by a plethora of parameters, is inherently challenging.
+7 Conclusion
+We introduced TimeFlow, a novel framework for longitudinal brain MR registration, specifically designed to address persistent challenges such as limited sequential data, the accuracy-temporal consistency trade-off, and the inability to predict future brain states.
+TimeFlow leverages a U-Net-based architecture with temporal condition embeddings to ensure temporally continuous and spatially smooth deformation fields, eliminating the need for explicit smoothness penalties.
+Unlike existing methods that require full sequential data, TimeFlow learns effectively from only two scans per patient, achieving both high registration accuracy and robust extrapolation to future timepoints.
+Extensive evaluations demonstrate TimeFlow’s superior performance over existing baselines in both interpolation and extrapolation tasks, as well as its ability to facilitate clinically meaningful analyses of brain aging and neurodegenerative diseases.
+By seamlessly bridging retrospective interpolation and prospective extrapolation, TimeFlow enables annotation-free analyses, representing a significant step toward data-driven models for tracking and predicting neuroanatomical changes in neurodegenerative diseases.

### v4

@@
-IEEEexample:BSTcontrol
+Temporal Conditioning for Longitudinal Brain MRI Registration and Aging Analysis
 Abstract
@@
 $\mathcal{L}_{\textrm{sim}}(I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L},I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}),$ (2)
-Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timpoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations.
+Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timepoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations.
 By freely sampling $\hat{t}$, TimeFlow effectively learns densely and continuously across the entire temporal axis.
@@
 To address this, we use “synthetic” triplets dynamically sampled from continuous time intervals.
-Specifically, by sampling intermediate timepoints $\hat{t}$ from $\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints.
+Specifically, by sampling intermediate timepoints from uniform distribution $\hat{t}\sim\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints.
 This continuous, densely sampled temporal supervision significantly enhances the network’s capacity to generalize and accurately model brain anatomy changes at any arbitrary timepoint, including unseen future scenarios, while using only two scans as inputs.
@@
 Data and Pre-processing
-We train and evaluate TimeFlow on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [4], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain.
-We employ the longitudinal processing from FreeSurfer v7.2 [49].
-First, all scans are processed individually, including skull stripping, normalization.
-Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [50].
+We train and evaluate TimeFlow on the publicly available Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [21], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both cognitively normal subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders characterized by progressive patterns of brain atrophy.
+All scans undergo standardized pre-processing using the longitudinal pipeline in FreeSurfer v7.2 [13].
+Each scan is first processed individually for skull stripping, normalization.
+Subsequently, all scans from the same subject are rigidly aligned via a groupwise rigid registration [41].
 FreeSurfer aggregates information across timepoints and refines the pre-processed outputs.
 To ensure meaningful longitudinal analysis, we select 134 subjects exhibiting substantial structural changes (median brain deformation >0.3>0.3mm).
-Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits.
-It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12.
-All scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels.
-The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping.
-Additionally, 11 subjects from the OASIS dataset [51] serve for zero-shot evaluation.
+The dataset is partitioned into training (74 subjects), validation (30 subjects), and testing (30 subjects) sets at the subject level prior to any processing, ensuring complete subject disjointness across splits and preventing information leakage.
+The splits are constructed to maintain balance across diagnosis, age, sex, and number of available visits.
+Each subject contributes to 2 to 12 longitudinal scans.
+All images are resampled to isotropic 1mm spacing and cropped to a uniform size of $160\times 160\times 192$ voxels.
+Intensity values are normalized to the range [0,1] based on the [0, 99.9] percentile without clipping.
+Additionally, 11 subjects from the OASIS dataset [24] are used for zero-shot evaluation to assess TimeFlow’s generalization to data from different scanners and imaging protocols.
 Figure 4: Visualization of extrapolation errors compared to ground truth future images. Coronal and axial views of two subjects are presented. Areas with high error are marked by red arrows.
@@
 Our model is implemented using Pytorch 2.1.2.
-and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per subject).
+and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per image pair).
 We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$.
@@
 […]
+[…]
+[…]
 Figure 6: Visualization of interpolation (Year 1.5) and extrapolation (Year 8) results from coronal and axial views of a subject’s brain.
@@
 Baselines
-To evaluate the registration performance, we compare TimeFlow with the following methods:
-  • NODER [17, 16] takes the complete image sequence as the model input and is thus aware of every time point.
-It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [52] to compute integration and adjoint gradients.
+To evaluate TimeFlow’s registration performance, we compare against the following methods:
+  • NODER [54, 3] takes the complete image sequence as input and is thus aware of every timepoint.
+It performs subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [8] to compute integration and adjoint gradients.
 It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
@@
 Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration.
-The number of integration steps of VXMdiff is also set to 7.
+The number of integration steps of VXMdiff is set to 7.
+  • SyN (ANTs) [1].
+We apply the symmetric diffeomorphic image pairwise registration algorithm with normalized cross-correlation (radius $r=3$).
+For interpolation and extrapolation, displacement fields are linearly scaled, analogous to VXM.
   • Sequence-based VoxelMorph (Seq-VXM) [53].
-Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point.
-Given that Seq-VXM performs multiple pairwise registrations by taking the full sequence as model input, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab.
+Seq-VXM processes the whole sequence and registers the baseline image to each intermediate timepoint through multiple pairwise registration.
+Since Seq-VXM uses the full image sequence as input, we treat it as the performance upper bound for the retrospective interpolation task (refer to Tab.
 1).
 All methods use Local Normalized Cross Correlation (LNCC) with radius $r=9$ (NODER) or $r=4$ (others) as image similarity loss.
-The loss function weighting factors in NODER follow their original suggestion.
+The loss function weighting factors for NODER follow the original implementation.
 Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [48, 47] with weight $1.5,0.5,1.0$.
@@
 ***:p<0.001, **:p<0.01, *:p<0.05
+Table 2: Dice Score on ADNI dataset. For each method, we report the mean Dice scores and standard deviation across all subjects and anatomical regions. Higher values indicate better alignment.
+Method Extrapolation Interpolation
+Seq-VXM - 81.17 $\pm$ 1.85
+SyN 72.85 $\pm$ 6.03 80.32 $\pm$ 2.76
+NODER 75.43 $\pm$ 4.43 80.43 $\pm$ 1.88
+VXM 75.78 $\pm$ 4.48 81.15 $\pm$ 3.14
+VXMdiff 75.35 $\pm$ 4.50 81.11 $\pm$ 2.75
+TimeFlow 76.58 $\pm$ 3.90 81.19 $\pm$ 2.88
+TimeFlowdiff 76.95 $\pm$ 3.80 81.10 $\pm$ 2.65
 Evaluation metrics
@@
 We first evaluate TimeFlow’s performance in prospective scenarios, specifically, its ability to extrapolate deformation fields to predict future brain scans.
-TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a surrogate metric to quantify the biological aging progression.
+TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a quantitative metric to quantify the biological aging progression.
 Given t, predict the future image
@@
 Note that Seq-VXM cannot carry out future predictions; thus, it is excluded from this comparison.
+In addition, the Dice score metric reported in Tables˜2 and 10 further confirms, in agreement with the intensity-based metrics, that TimeFlow achieves better anatomical region alignment.
 Qualitative extrapolation results are visualized in Figure˜6 (Year 8) and Figure˜4, illustrating TimeFlow’s accuracy by comparing to the ground truth images.
@@
 Given future scan, assess the brain aging
-As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis.
+As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis (i.e., requiring no anatomical segmentation or surface reconstructions for quantification).
 We compare the brain aging rates across diagnostic groups, i.e., Controls (CN), MCI, and Dementia, by predicting the optimal time parameter ($t_{\textrm{ext}}$) corresponding to the observed future MRI scans.
@@
 Indeed, the boxplot of Figure˜7 (Left) clearly shows that Dementia patients exhibit the highest progression rate, followed by MCI subjects, while the control group maintains progression rates around 1.
-The analysis validates that our annotation-free surrogate brain aging rate metric is able to separate disease groups.
+The analysis validates that our annotation-free brain aging rate estimate is able to separate disease groups.
+Figure 9: Stability analysis across interpolation and extrapolation regimes.
+Left: Mean Absolute Error (MAE, in units of $\times 10^{-2}$) of TimeFlowdiff as a function of time $t$.
+The solid line shows the Gaussian kernel-smoothed mean, and the shaded region represents $\pm 1$ standard deviation across subjects.
+Right: Non-diffeomorphic Volume (NDV, logarithmic scale) of TimeFlowdiff as a function of time $t$.
+Hollow dots indicate NDV =0=0 (perfectly diffeomorphic).
+The vertical dashed line at $t=1$ separates interpolation ($0<t<1$) from extrapolation ($t>1$).
+The x-axis uses a piecewise linear scale: the left segment spans $t\in[0,1]$ and the right segment spans $t\in[1,10]$.
+Figure 10: Region-wise Dice score comparison for extrapolation on ADNI dataset.
+Each boxplot shows the distribution of Dice scores across subjects for major anatomical structures.
+TimeFlow variants achieve consistently higher or comparable Dice scores across all regions.
+Figure 11: Region-wise Dice score comparison for interpolation on ADNI dataset.
+Each boxplot shows the distribution of Dice scores across subjects for major anatomical structures.
+All methods achieve high Dice scores ($80\%$), with TimeFlow variants performing comparably to the upper-bound Seq-VXM.
 […]
@@
 Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without extra spatial smoothness regularization.
+Region-wise segmentation-based evaluation (Tables˜2 and 11) validates the superior performance of TimeFlow and TimeFlowdiff.
 Qualitative visualizations in Figure˜6 (Year 1.5) illustrate the accuracy of TimeFlow’s interpolations, substantially reducing alignment errors compared to initial misalignments.
-Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a surrogate brain progression state marker.
+Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a quantitative index of brain progression state marker.
 We extract trajectory segments with the same diagnoses from each subject, and then compute a linear relationship between the estimated brain state ($t$) and chronological age (years since baseline of the segment).
 The median linear lines of the groups (CN, MCI, Dementia) in Figure˜7 (Right) clearly illustrate distinct progression patterns, with Dementia exhibiting the fastest structural changes, MCI intermediate, and CN the slowest.
+5.3 Clinical Validation, Interpretability and Robustness
+Correlation with Cognitive Decline
+To further validate the clinical relevance of TimeFlow’s learned trajectories, we assess the association between the model’s estimated brain state ($t$) and cognitive performance using Repeated Measures Correlation (rmcorr).
+This analysis reveals a significant negative correlation (r=0.46r=-0.46, 95% CI: [0.51,0.40][-0.51,-0.40]) between the inferred brain state ($t$) and Mini-Mental State Examination (MMSE) scores, indicating that greater pathological progression is associated with declining cognitive function.
+Importantly, this correlation emerges from trajectories constructed purely from anatomical features of the input image pairs, without any supervision from cognitive scores or diagnostic labels during training.
+This finding demonstrates that TimeFlow’s unsupervised deformation modeling effectively captures clinically meaningful, subject-specific disease progression patterns.
+Distinction from diagnostic classification
+Unlike supervised diagnostic classification models that predict discrete disease labels, TimeFlow operates as an unsupervised and disease-agnostic framework designed to capture continuous spatiotemporal structural changes.
+The model receives no explicit disease labels during training.
+Instead, it learns to model anatomical trajectories from image pairs alone.
+The strong alignment between TimeFlow’s derived progression metrics and diagnostic groups (Figure˜7) emerges naturally from the learned deformation fields, confirming that the framework effectively quantifies pathological deviations without disease-specific priors.
+This unsupervised nature enables TimeFlow to provide granular, interpretable progression quantification that complements categorical diagnoses.
+Moreover, by generating predicted future brain images, TimeFlow offers clinicians direct visual intuition: "if the subject continues progressing at the current rate, what will their brain look like in, e.g., 2 years?" This capability bridges quantitative biomarkers with interpretable visual assessment, supporting more informed clinical decision-making.
+Stability of extrapolation
+To assess TimeFlow’s robustness across both interpolation and long-range extrapolation, we analyze registration accuracy (MAE) and deformation regularity (NDV) as functions of time $t$ in Figure˜9.
+For interpolation ($0<t<1$), TimeFlow achieves consistently low MAE (below $8\times 10^{-2}$) and diffeomorphic deformations (NDV =0=0).
+For extrapolation ($t>1$), MAE increases monotonically with prediction distance, reflecting the inherent uncertainty in long-term forecasting.
+However, for near-to-mid-term predictions ($1<t\leq 5$), the error remains stable below $10\times 10^{-2}$, comparable to the mean extrapolation MAE reported in Table˜1.
+Crucially, NDV remains negligible ($0.01\%$ of total brain volume) for $t\leq 5$, indicating minimal folding artifacts.
+These results confirm that TimeFlow generates anatomically plausible deformations and maintains stability for clinically relevant prediction horizons.
 5.4 Ablation Study
@@
 Specifically, we compute the norm of the temporal displacement derivatives within the brain regions and then take the mean over all brain voxels as the final metric.
-As shown in table˜2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy.
+As shown in table˜3, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy.
 Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to generate plausible extrapolated deformation fields for future image estimation.
 These ablation results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields, achieving an optimal balance between registration accuracy and smoothness without explicit smoothness penalties.
+5.5 Runtime analysis
+We compare computational efficiency across all methods.
+Optimization-based methods exhibit substantially longer runtimes: SyN requires approximately 5 minutes per image pair on CPU, while NODER performs subject-specific optimization requiring 1-2 minutes per subject on GPU.
+In contrast, learning-based methods achieve efficient amortized inference.
+TimeFlow and TimeFlowdiff complete registration in 5-7 seconds per subject, VXM and VXMdiff in 1-2 seconds, and Seq-VXM in 3-5 seconds (all on GPU).
+Beyond runtime, memory requirements differ significantly for long sequences: NODER and Seq-VXM require over 16 GB GPU memory for sequences exceeding 5 timepoints, whereas TimeFlow operates efficiently with less than 8 GB.
 6 Discussion
@@
 Discrepancies between predicted and observed results could then be attributed to treatment effects, offering a novel and efficient approach for evaluating therapeutic efficacy in clinical settings.
+Distinction from Continuous-time and Generative Frameworks
+Although NeuralODE-based approaches like NODER [3, 54] share the capability of continuous trajectory modeling, they differ fundamentally from TimeFlow in both computational efficiency and modeling flexibility.
+NODER parametrizes a time-varying velocity field and require computationally expensive numerical integration of ordinary differential equations (ODEs).
+In contrast, TimeFlow learns a direct conditional mapping $\phi_{t}=f_{\theta}(I_{0},I_{L},t)$, bypassing explicit ODE solvers to enable computationally efficient, amortized inference.
+Furthermore, NODER relies on rigid chronological timestamps from the actual image acquisition times, constraining the trajectory to strictly follow the observed temporal sequence.
+TimeFlow, however, employs synthetic triplet sampling during training, which provides greater modeling flexibility and robustness to irregular temporal intervals commonly encountered in longitudinal datasets.
+This design enables TimeFlow to perform reliable inter-/extra-polation even when scan intervals vary substantially.
+Regarding generative approaches, while TimeFlow shares architectural similarities with diffusion-based generative models (e.g., U-Net with time embeddings), the objectives differ fundamentally.
+Diffusion models synthesize image intensities via stochastic denoising processes, generating visually realistic images but operating largely as black boxes with limited interpretability.
+In contrast, TimeFlow predicts deterministic, biophysically plausible deformation fields that enable direct quantification and interpretation of morphological changes.
+This distinction makes TimeFlow particularly suitable for clinical applications requiring explainable predictions and precise anatomical correspondence tracking.
 Bidirectional extrapolation
@@
 Backward extrapolation, i.e., traveling back in time and visualizing the brain at earlier, unobserved ages, opens several avenues: (i) reconstructing a pre-symptomatic baseline for patients before neurodegeneration; (ii) generating counterfactual controls that share a subject’s anatomy but simulate a healthy aging trajectory.
+Generalization and Future Outlook
+The principles underlying TimeFlow extend well beyond the context of brain aging and neurodegeneration.
+The framework can generalize to other populations, such as the opposite end of the lifespan: the infant brain, which exhibits rapid structural changes during early development.
+Moreover, the proposed inter-/extrapolation consistency constraints are inherently modality- and organ-agnostic, opening the possibility for application to other imaging modalities (e.g., T2-weighted, FLAIR) or even other organ systems.
+For example, by adapting the time embedding to capture the cyclic motion of the heart, TimeFlow could potentially model cardiac dynamics throughout the cardiac cycle.
+We acknowledge, however, that pure registration has limits: pathologies involving distinct intensity changes or new tissue appearance, such as tumors, inflammatory lesions, or acute stroke, require both geometric modeling and intensity synthesis.
+In such cases, combining TimeFlow’s precise deformation fields with generative models for intensity prediction could provide a more comprehensive approach to disease progression modeling.
 Limitation and failure cases